OpenStudy (anonymous):

Poker hands usually consist of 5 cards. What is the probability of being dealt a poker hand consisting of ace, king, queen, jack and ten all in the same suit? (royal flush). URGENT HELP

OpenStudy (anonymous):

20/2 598 960 4/2 598 960 1/2 598 960 5/2 598 960 which option would it be?

OpenStudy (anonymous):

anyidea?

OpenStudy (anonymous):

can u help me with this question: You draw a card from a deck of playing cards. What is the probability that it is a heart or a spade?

OpenStudy (anonymous):

P ( heart or spade ) = P ( heart) + P (spade) = ( number of heart + spade ) / 52

OpenStudy (anonymous):

so 1/2

OpenStudy (anonymous):

yes ( it's easy to check when you look at your deck of card)

OpenStudy (anonymous):

and did u get the poker question? like which option it would be

OpenStudy (anonymous):

= 4/52 * 4/51 * 4/50 * 4/49 * 4/48

OpenStudy (anonymous):

Could you try to plug them into your calculator?

OpenStudy (anonymous):

i did i get a weird number

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@mertsj can u help?

OpenStudy (anonymous):

Oh, divided by 2 598 960

OpenStudy (mertsj):

5/52x4/51x3/50x2/49x1/48=1/2598960

Directrix (directrix):

Continuing form above, the probability is 0.00000153908.